OpenStudy (anonymous):

A policeman on a motorcycle is pursuing a car that is speeding at 115 kmper hour. The policeman is 6 km behind the car and is traveling 130 kmper hour. How long will it be before the policeman overtakes the car? answer it in equation writing... plsss....

OpenStudy (hoblos):

130x +6 = 115x 15x = 6 x=6/15 h = 24min

OpenStudy (anonymous):

130-115 = 15km. 15kmX(t)hr = 6km so t = 6/15 hours. But there are 60 mins in an hour, so multiply this by 60 to get t=360/15 minutes = 24mins